OpenStudy (anonymous):

If the range of the function f(x) = 4x – 3 is {11.4, 15, 17, 29}, what is its domain? A) {1.5, 2.8, 5.5, 7} B) {2.3, 4, 8, 11} C) {7, 11, 13.2, 14} D) {3.4, 5, 8, 9.3} E) {3.6, 4.5, 5, 8}

OpenStudy (anonymous):

Lol are you doing plato?

OpenStudy (anonymous):

yus @kathyzhu23

OpenStudy (dan815):

f(x) = 4x – 3 is {11.4, 15, 17, 29}<----- These are the f(x) values You can work backwards to see what the x values must have been, the x values are your domain. 11.4=4x-3, x=? 15=4x-3, x=? 17=4x-3,x=? 29=4x-3,x=?

OpenStudy (anonymous):

its E ?